# Smart Parking with Pixel-Wise ROI Selection for Vehicle Detection Using YOLOv8, YOLOv9, YOLOv10, and YOLOv11

### v1

@@ 1 Introduction
  • A comparison of inference times for four different YOLO versions across six different hardware platforms with varying computational capacities. Additionally, we measured the latency in a standardized GPU environment (NVIDIA A100), which is not found on the literature.
@@ 4.2 Inference Model
A lot of classifiers could be used and compared, as, there is a wide range of possible models in terms of object detection, such as shown by Table 1.Sapkota et al. [2024][31] made an analysis of the advancements from traditional approaches such as classic machine learning classifiers, followed by the emergence of CNN’s, Region-based CNN, and the YOLO series. As stated by other works superior results of YOLO [32][33] and initial tests not showing promising results with models such as Masrk R-CNN and EfficientDet, we selected different recent flavors of YOLO in this study, with versions 8, 9, 10, and 11, selecting the largest and lightest model of each version (eg. YOLOv8x and YOLOv8n).
@@
YOLOv8 [40] uses an anchor-free model with a decoupled head, sigmoid function for activation, and softmax function for class probabilities, using the backbone as a variant of CSPNet with the named cross-stage partial bottleneck with two convolutions (C2f) module, combining high-level features with contextual information to improve performance [41]. After going throught the CNN, the FPN is replaced by a PAN.
@@ 5.1 Model Performance
It is possible to observe that despite the model chosen, the post-processed method showed higher balanced accuracy, with significant gains such as over 28 percent points at YOLOv8n. It was possible no notice that, as expected, the lightest models performs worse than the heaviest, but the range of this variation is different according to the model evaluated. For example, specially when analysing YOLOv9t and YOLOv9e with the post-processed method, the balanced accuracy goes from 86.90% to 99.68%, indicating a range of around 13 percent points. This variation was not strongly observed at the other YOLO versions. YOLOv10, in particular, showed to be an stable model for both ROI methods, despite not having the highest performance. YOLOv11 presented a high performance specially at the extra large version, however still lower than YOLOv9, which was the better model in terms of accuracy, balanced accuracy and $F_{1}$-score. To continue the analysis, we investigated the confusion matrices for the worst and best possible result for each ROI method at Figures 7 and 8.
@@ Acknowledgments
The datasets generated and/or analysed during the current study are not available due to the sensitive nature of the research, however codes and materials that support the findings of this study are available upon reasonable request. All licence plates that could appear in this work are already unidentified at the base picture and also were blurred.

### v2

@@ 1 Introduction
  • comparison of inference times for four different YOLO versions across six different hardware platforms with varying computational capacities. Additionally, we measured the latency in a standardized GPU environment (NVIDIA A100), which is not found on the literature.
@@ 4.2 Inference Model
A lot of classifiers could be used and compared, as, there is a wide range of possible models in terms of object detection, such as shown by Table 1. Sapkota et al. [2024][31] made an analysis of the advancements from traditional approaches such as classic machine learning classifiers, followed by the emergence of CNN’s, Region-based CNN, and the YOLO series. As stated by other works superior results of YOLO [32][33] and initial tests not showing promising results with models such as Mask R-CNN and EfficientDet, we selected different recent flavors of YOLO in this study, with versions 8, 9, 10, and 11, selecting the largest and lightest model of each version (eg. YOLOv8x and YOLOv8n).
@@
YOLOv8 [40] uses an anchor-free model with a decoupled head, sigmoid function for activation, and softmax function for class probabilities, using the backbone as a variant of CSPNet with the named cross-stage partial bottleneck with two convolutions (C2f) module, combining high-level features with contextual information to improve performance [41]. After going through the CNN, the FPN is replaced by a PAN.
@@ 5.1 Model Performance
It is possible to observe that despite the model chosen, the post-processed method showed higher balanced accuracy, with significant gains such as over 28 percent points at YOLOv8n. It was possible to notice that, as expected, the lightest models performs worse than the heaviest, but the range of this variation is different according to the model evaluated. For example, specially when analysing YOLOv9t and YOLOv9e with the post-processed method, the balanced accuracy goes from 86.90% to 99.68%, indicating a range of around 13 percent points. This variation was not strongly observed at the other YOLO versions. YOLOv10, in particular, showed to be an stable model for both ROI methods, despite not having the highest performance. YOLOv11 presented a high performance specially at the extra large version, however still lower than YOLOv9, which was the better model in terms of accuracy, balanced accuracy and $F_{1}$-score. To continue the analysis, we investigated the confusion matrices for the worst and best possible result for each ROI method at Figures 7 and 8.
@@ Acknowledgments
The datasets generated and/or analyzed during the current study are not available due to the sensitive nature of the research, however codes and materials that support the findings of this study are available upon reasonable request. All license plates that could appear in this work are already unidentified at the base picture and also were blurred.